C2H4(g) + 3O2(g) à 2CO2(g) + 2H2O(g)

If 2.7mol of ethane is reacted with 3.6mol of oxygen

a)Identify the limiting and excess reactants

b)Find the amount of both products produced

c)Find how much of my excess reactant is left over at the end of the reaction
Answer #1

a)
Balanced chemical equation is:
C2H4 + 3 O2 ---> 2 CO2 + 2 H2O


1 mol of C2H4 reacts with 3 mol of O2
for 2.7 mol of C2H4, 8.1 mol of O2 is required
But we have 3.6 mol of O2

so, O2 is limiting reagent
Answer: O2

b)
According to balanced equation
mol of CO2 formed = (2/3)* moles of O2
= (2/3)*3.6
= 2.4 mol

According to balanced equation
mol of H2O formed = (2/3)* moles of O2
= (2/3)*3.6
= 2.4 mol

Answer:
mol of CO2 formed = 2.4 mol
mol of H2O formed = 2.4 mol

c)
According to balanced equation
mol of C2H4 reacted = (1/3)* moles of O2
= (1/3)*3.6
= 1.2 mol
mol of C2H4 remaining = mol initially present - mol reacted
mol of C2H4 remaining = 2.7 - 1.2
mol of C2H4 remaining = 1.5 mol
Answer: 1.5 mol